OpenStudy (anonymous):

if f(x) approximate g(x) within 1/3000 for 10

OpenStudy (anonymous):

you always come up with the damndest questions but lets say \(f(x)=0\) and \(g(x)=\frac{1}{3000}\) which is as far as it can be then \[\int_{10}^{12}f(x)dx=0\] and \[\int_{10}^{12}\frac{1}{3000}=\frac{2}{3000}\]

OpenStudy (anonymous):

a more sophisticated way to answer would be to say that \[\int f -\int g =\int (f-g)\leq \int \frac{1}{3000}=\frac{2}{3000}\]

OpenStudy (anonymous):

an even more sophisticated way to answer would be to use absolute values, but enough already

OpenStudy (anonymous):

did you get an answer to that series with the squares? that bedevilled me for a while

OpenStudy (anonymous):

the series with squares? which one..i don't even remember..lol

OpenStudy (anonymous):

\[x+2^2x^2+3^2x^3+4^2x^4+...\]

OpenStudy (anonymous):

or maybe it started with 1, i don't remember but i could not figure out how to get that square there

OpenStudy (anonymous):

i think the series is e^x expansion and minus first two term....lol right?

OpenStudy (anonymous):

yeah right your questions vary from more or less immediately obvious to wtf with very little in between

OpenStudy (anonymous):

maybe a bi-polar professor?

OpenStudy (anonymous):

i don't know ..i'm taking an online calculus...it's really frustrate me